Mix a solution that is 30% alcohol with a solution that is 80% alcohol to make 800mL of a soultion that is 60% alcoho. How much of each solution should you use?
Answer by ankor@dixie-net.com(22740)

Mix a solution that is 30% alcohol with a solution that is 80% alcohol to make 800mL of a solution that is 60% alcohol. How much of each solution should you use?
:
Let x = amt of 30% solution, since the total amt is 800 ml, then
(800-x) = amt of 80% solution.
:
Convert % to decimal and write an equation:
:
.30(x) + .80(800-x) = .60(800)
:
.3x + 640 - .8x = 480
.3x - .8x = 480 - 640
-.5x = -160
x = -160/-.5
x = + 320 ml of the 30% solution
:
Find the amt of the 80% solution required; 800 - 320 = 480 ml
:
Check: .3(320) + .8(480) = .6(800)